# Development of a time correction algorithm for a precise synchronization of a free-running Rubidium atomic clock with the GPS Time

### v1

@@ 2.2.2 Validation of the method with simulations
Before evaluating the performance of our timing system when integrating the correction algorithm, the method was validated on simulated signals [27] in order to isolate the effect and performance of the correction from any measurement effect.
@@
Three types of signals were considered: a perfect clock to be used as a reference to evaluate the performance, a free-running Rubidium clock and a GPS time signal, as measured by the Septentrio receiver. The quadratic drift was not included because it is deterministic and therefore does not require further study for being corrected. At first order, the clock signal can be modeled by white noise ($WN$) in both phase and frequency as well as a random walk ($RW$) noise in frequency. Based on the characterization of the Rb clock, the phase and frequency flicker noises can be neglected for this purpose. Indeed, the characterization of our Rubidium clock in Figure 2 showed that the frequency flicker noise had a negligible impact on the OASD. Furthermore, the phase white and flicker noises have a similar impact on the standard OASD and cannot be distinguished here. We chose to ignore the phase flicker noise as it is less straightforward to simulate and it should not impact the long term random walk that we want to correct. The GPS Time can be modeled as pure phase white noise. The corresponding OASD as a function of the averaging time $\tau$ can be modeled [28, 29, 30] by:
$OASD(\tau)\cong A_{WNp}\times\tau^{-1}+A_{WNf}\times\tau^{-1/2}+A_{RWf}\times\tau^{+1/2}.$ (2.3)
The amplitudes $A$ of these main frequency and phase noises were determined through fitting this model (Eq. 2.3) to the OASD of the data when characterizing our equipment (see Figure 2) and found to be:
$A_{WNf}$ $=$ $7\times 10^{-12}~{}s^{1/2},$ (2.4)
@@
for the free-running Rb clock and for the GPS Time:
$A_{WNf}$ $=$ $0~{}s^{1/2},$ (2.5)
@@
In Figure 7, the uncorrected simulated signals of the GPS and the clock are reported in dotted symbols for comparison. The increase of the clock’s OASD after $\tau=10^{4}$ s due to the random walk is clearly visible. One can see that the OASD of the corrected signals (starred symbols) do eliminate the random walk at longer terms which indicates a success of the correction method (quadratic). Moreover, one can determine that the ideal length $\Delta t$ of the correction time windows lies around $3\times 10^{4}$ s which corresponds logically to the intersection of the free-running Rb clock and GPS Time OASD curves. Indeed, the red curve with a time window of 28800 s shows an ideal combination of the short-term stability of the clock and the absence of random walk at longer scales. On the opposite, the yellow (shorter time window) and light blue (longer time window) curves show respectively a degradation of the short term performance and a remaining random walk component in the region between $\tau=10^{4}$ s and the time window length (here 240000 s).
@@
Finally, it is important to note that although this validates the methods for application on data, those are simplified simulations, in particular because only the main noise types are taken into account. As a result, we do expect differences of performance of the correction on real data. It is also possible that the optimal time window for the correction is slightly different for real data because the simulations are not exact representation of data. Two main differences can be noted: the absence of frequency drift and flicker noises in the simulated Rubidium signal and the fact that we assume a perfect signal to compare the Rubidium signal to when evaluating the OASD.
@@ 2.2.3 Implementation on data
To check the impact of the correction we compare the Rubidium clock signal to that of another more stable clock, like a Passive Hydrogen Maser. The PHM signal plays the role of the perfect signal used for the simulations, while obviously not being perfect. This first difference is to take into account while comparing performances on simulated data to performance on experimental data. In the following, we will also quantify the stability of the Rubidium signal using the OASD of a series of frequency ratios (according to equation (10) of [31]) between this signal and the $5$ MHz generated by the PHM. Measuring this ratio frequently, once per second for instance, would allow to also evaluate the very short term stability of the corrected signal which is not possible with the Septentrio measurements that are integrated over $16$ minutes. We use the frequency counter to provide a measurement per second of the Rubidium clock 5 MHz frequency $f_{Rb}^{i}$ taking the PHM 5 MHz generated signal as a frequency reference $f_{ref}$. We then perform a simultaneous correction of the Rubidium - GPS Time, as measured by the Septentrio receiver, and of this frequency ratio $f^{i}=f_{Rb}/f_{ref}$ series. Comparing the OASD of the corrected frequency series to the uncorrected one, one can quantify the short term stability (below $16$ minutes) after correction while making sure that the random walk was corrected. We can also use this comparison to optimize the value of $\Delta t$ in order to achieve the lowest Allan Standard Deviation possible at all averaging time windows.
@@ 3 Results
In this Section, we present the results of the correction of the Rubidium clock time stamps obtained for simultaneous measurements of $\sim 50$ days with the Septentrio receiver and the frequency counter with PHM $5$ MHz signal as a frequency reference. The frequency measurements are divided by the expected value to obtain a series of Rb/PHM frequency ratios. The OASD of such a frequency series is shown in Figure 2. Note that the statistical uncertainty on the estimated OASD, due to the limited number of samples per averaging time, are included as error bars for both curves (Rb and GPS) but they are too small to be visible. Indeed for the Rb vs PHM OASD, the statistical uncertainty is at the permil level. Up to an averaging time of around $10^{4}$ s, the stability is limited only by the phase white noise and then by the frequency white noise. After that, the OASD first increases as $\tau^{1/2}$ which is characteristic of the frequency random walk. From $\tau\approx 5\times 10^{5}$ s, the OASD increases proportionally to $\tau$. This is characteristic of a deterministic frequency drift which can be easily characterized and corrected for contrary to the frequency random walk. In comparison, the OASD of the difference between GPS Time and UTC(OP) that we receive from the SYRTE laboratory via White Rabbit network, is only limited by a phase white noise at least up to an averaging time of $5\times 10^{5}$ s: the OASD keeps decreasing with the averaging time. At low averaging times, the GPS stability is worse than that of the Rb because of this phase white noise: the GPS OASD is of around $3\times 10^{-12}$ at $960$ s compared to around $7\times 10^{-13}$ OASD for the Rubidium clock. However, at around $10^{4}$ s, the stability of the Rb signal becomes worse compared to GPS Time because of the frequency random walk and drift of the Rubidium clock.
@@
In this paper, we used only the GPS satellites with an elevation angle (angle between line of sight and horizontal direction) larger than $15^{\circ}$ to extract the Rubidium time residuals distribution. During the whole data-taking period, for each data point, the Septentrio receiver was able to track an average of $6.5$ GPS satellites and at least $4$ GPS satellites for each data point. To obtain the Rubidium vs GPS Time difference, we take the mean value of the differences between the Rubidium clock and each GPS satellite tracked in the same integration time window of the Septentrio receiver. The obtained time difference is shown in Figure 10. The time difference shown here have already been corrected for the deterministic drift discussed before as this can be easily monitored and corrected for contrarily to the random walk. To correct for both this frequency drift, we performed a quadratic fit of the first few days of Septentrio measurements and subtracted the results to the Septentrio data. The frequency drift was measured at around $10^{-18}~{}s^{-1}$ which would induce a drift of $100$ ns after less than four days. Note that the correction of the frequency deterministic drift could be actualized regularly with new fits of the time difference over a long enough period (a few days of data) but it was not found to be necessary for our example with 50 days of data-taking. In the following, the same initial correction of the frequency drift was applied to the Rb vs PHM data. The correction coefficients will be extracted from the residual time difference. Before correction, we see that after a few days of data-taking, the Rb clock can drift away from the GPS Time by more than a hundred nanoseconds because of the random walk noise.
@@
Figure 10: Time difference between the Rubidium clock and GPS Time as measured by the Septentrio receiver. The data are already corrected for the deterministic drift of the Rubidium clock. The large variations are caused by the frequency random walk of the Rubidium clock.
@@ 3.1 Offline correction
Figure 11 shows the Allan Standard Deviation of the Rubidium vs PHM data. Note that a relative resolution of $10^{-11}$ was chosen for the frequency measurement with the frequency counter. This is lower than the Phase White Noise of the free-running Rubidium clock so it does not impact significantly the Allan Standard Deviation. The blue curve shows the result for series corrected only for the deterministic drift of the Rubidium clock, by subtracting the expected time distribution of the series caused by this drift. Note that by correcting the deterministic drift, we also partially correct for the frequency random walk such that the OASD decreases with the averaging times for $\tau>10^{6}$ s. The other colored curves show the results for the series corrected offline, with different width of the correction time window. Here, we use quadratic fits of the Septentrio data (so $a_{k}\neq 0$ a priori). The frequency series were integrated to obtain time series to which we subtracted the fit results. The shortest time window ($2880$ s) corresponds to approximately $3$ Septentrio $16$ minutes epochs. The medium ($10560$ s) and largest ($240,000$ s) correspond respectively to $11$ and $250$ Septentrio data points.
@@
Figure 11: Overlapping Allan Standard Deviation of the Rb/PHM frequency ratios series after the deterministic drift correction (in blue) and after the correction with a correction time window of $2880$ s (orange), $10560$ s (green) and $240,000$ s (red). The best stability at both short and long averaging times is obtained for the medium time window ($10560$ s3absent3$\approx$ 3≈ 3 hours).
@@
One sees that with the medium time window compared to the two others, we obtain the best stability at all averaging times. At lower averaging times, the performance is very similar to the uncorrected time series. At higher averaging times, the Allan Standard Deviation is much better than the uncorrected series and is comparable to the one obtained for the shortest correction time window. This illustrates the fact that both the $2880$ s and $10560$ s windows are able to correct very well the frequency random walk ($\tau^{1/2}$ component of the ASD) of the uncorrected time series. However, with the shortest correction time window, the short term stability of the time series is degraded compared to the uncorrected series: the value of the ASD at $1$ s increases by a factor $\sim 1.5$. In this scenario, the corrected Rubidium time signal gets very close to GPS Time which is known to have a higher phase White Noise. Finally, the longest correction time window leads to a similar stability as the shortest one at long term, and even poorer stability at $\tau\in[10^{4},10^{5}]$ s.
Figure 12 shows the Rubidium vs GPS Time difference after the offline correction. The shorter the correction time window, the better. However, with the medium length time window, we still get time residuals lower than $5$ ns over the whole data-taking period, which is well below the requirements of HK. With the longest correction time window, jumps of a few tens of nanoseconds are introduced in the time residuals. This explains the overall higher ASD: the stability of the signal is limited by those jumps. These jumps can be understood by looking at the fit of the Septentrio data in this scenario in Figure 13. The time scale of the variations in the data to fit is too small compared to the $240,000$ s time window. In consequence, the fitted tendency from one piece to another is very different, and the fitted piece-wise polynomial is not continuous.
@@
Figure 12: Time difference between the Rubidium clock and GPS Time after the offline correction. Three different correction time windows have been tested: $2800$ s (orange), $10560$ s (green) and $240,000$ s (red). These residuals can be compared to the residuals before correction that were shown in Figure 10.
Figure 13: Time difference between the Rubidium clock and GPS Time. The red portions show the results of the polynomial fit over consecutive time windows of $240,000$ s. The fit sometimes fail to represent the shorter time scales variations of the measured data. The poor fit quality can then lead to introducing jumps in the corrected time signal.
@@
With the offline version of the corrections, we thus obtain a very good synchronization to GPS Time at the level of a few nanoseconds with the $10560$ s time window. However, this version of the correction cannot be applied in real time. In the following paragraphs, we show the results for the online version of the correction that can be applied in real time to correct the time stamps of events in physics experiments.
@@ 3.2 Online correction
Figure 14 shows the Allan Standard Deviation of the uncorrected (blue) and online corrected (other colors) Rubidium vs PHM frequency times series. The same three time window intervals as in the offline correction scenario are considered. The top panel shows the results using quadratic fits of the Septentrio data and the bottom panel shows the results with linear fits. For the shortest and medium correction time windows, the linear fits lead to better performance with a lower OASD at low averaging times. At $1$ s, the OASD with the shortest (medium) correction time window is reduced by a factor $4$ (resp. $\sim 1.5$).
@@
Figure 14: Overlapping Allan Standard Deviation of the Rb/PHM frequency ratios series after the deterministic drift correction (in blue) and after the online correction with a correction time window of $2880$ s (orange), $10560$ s (green) and $240,000$ s (red). The data were fitted with quadratic (top) or linear (bottom) functions of time. A better stability, similar to the offline correction, can be obtained using linear fits.
@@
This behavior is very understandable looking at the number of degrees of freedom (number of data points - number of free parameters) in our fits. For the shortest time windows, the number of degrees of freedom is relatively low (00 and $8$) in case of quadratic fits so we risk over-fitting to the past data in order to correct the present data. This number of degrees of freedom is less relevant in the offline correction as the fit is performed on the same data as the correction (the over-fitting is not a problem here). Lowering the number of free parameters is one way of increasing the degrees of freedom hence allowing the fit to better generalize to the present data. Another way to increase the number of degrees of freedom is to increase the number of data points in the fit. For the longest time window, there are $247$ degrees of freedom in the quadratic fit so we lower the risk of over-fitting. On the contrary, in that case, quadratic fits lead to a slightly better correction of the random walk that limits the stability only up to $\tau\approx 8\times 10^{4}$ s whereas with linear fits, it limits the stability up to $\approx 2\times 10^{5}$ s. Here, Figure 14 shows a clear degradation of the stability after correction for averaging times lower than the correction window’s length. This is a known effect from linear servo loop theories and periodic perturbations of oscillators [32] and it could be attenuated by scaling down the correction: instead of subtracting the result of the fit, we could subtract only a fraction of it. Additionally, degradation of short term stability because of over-fitting on data from the past is, also visible in the OASD of the Rb vs GPS time difference after correction shown in Figure 15. This plot also illustrates that the corrected series stability, compared to GPS Time, is not limited by any frequency random walk at least up to an averaging time of $2\times 10^{6}$ s and with a correction time window short enough. Indeed, in that case, the OASD keeps decreasing with increasing averaging time.
Figure 15: Overlapping Allan Standard Deviation of the Rb vs GPS Time residuals series after the deterministic drift correction (in blue) and after the online correction with a correction time window of $2880$ s (orange), $10560$ s (green) and $240,000$ s (red). Note that before the deterministic drift correction, the OASD of this signal is the combination of the GPS Time OASD at low $\tau$ and the Rubidium OASD at high $\tau$, i.e: the combination of the blue and orange curves of Figure 2. The deterministic drift correction slightly smooths the residuals so that the OASD becomes generally lower and the frequency drift and random walk at high $\tau$ disappears, hence the decreasing OASD in the blue curve at very high $\tau$. The time residuals were fitted with quadratic functions of time. The increase of OASD at $1$ s averaging time with decreasing correction time window is consistent with what is observed in the Rb/PHM frequency ratios series after online correction. The less degrees of freedom in the fit, the more we risk over-fitting on past data and lowering the short term stability of the signal.
@@
Regarding the stability of the corrected Rubidium clock, using linear fits, the conclusions are the same as for the offline correction. The lowest Allan Standard Deviation, for all averaging times, is achieved with the medium width correction time window. With the shortest time window, the short term stability is degraded, whereas it is the long term stability that is degraded (compared to the other corrected scenarios) with the longest correction time window. Note that, contrary to the offline correction, the online correction with very long time windows does not deteriorate the short term stability. This is due to the use of "overlapping" windows of Rb vs GPS data. Between two consecutive fits, there is only one data point out of the 250 used that changes (the oldest one from the previous fit is replaced by the newest point). The fit parameters cannot change too abruptly from one fit to another so the resulting distributions are smooth.
@@
If the correction time window is too wide, we cannot correct as well the frequency random walk of the free-running Rubidium: the risk is that the Rubidium time signal locally drifts too far away from the GPS Time. This can be observed in the corrected Rubidium against GPS Time in Figure 16 where the maximum difference reaches $\sim 60$ ns (or $\sim 25$ ns with quadratic fits) with the $240,000$ s correction time window. With the $10560$ s correction time window, the differences stay in the ±$\pm 5$ ns range. Once again, one can see the reduction of the white noise when using linear instead of quadratic fits for the $2880$ s correction time window scenario: the residuals are contained in a ±$\pm 5$ ns range with linear fits instead of ±$\pm 12$ ns with quadratic fits. Before correction, as the reader saw in Figure 10, the free-running Rubidium clock can drift by around $100$ ns in $10$ days which means that HK’s requirement for the synchronization with UTC is not met. After online correction with the longest time window tested, the corrected Rubidium time stamps drift by around $60$ ns in a few days because of remaining random walk noise. Even though during the $50$ days data-taking period the time residuals with respect to GPS Time does not exceed $100$ ns, it is not possible to safely claim that the Rubidium clock drift will not exceed HK’s requirement of $100$ ns if we use the $240,000$ s correction time window. With shorter time windows, this drift seems to be dominated by white noise and is thus contained in a range of a few nanoseconds.

### v2

@@ 2.2.2 Validation of the method with simulations
Before evaluating the performance of our timing system when integrating the correction algorithm, the method was validated on simulated signals [27] in order to isolate the effect and performance of the correction from any measurement issues.
@@
Two types of signal comparisons to a perfect reference were simulated: a free running Rubidium clock and a GPS time signal, as measured by the Septentrio receiver. The quadratic time drift due to the Rubidium clock frequency drift was not included because it is deterministic and is therefore not challenging to correct. At first order, the clock signal can be modeled by white noise ($WN$) in both phase and frequency as well as a random walk ($RW$) noise in frequency. Based on the characterization of the Rubidium clock, the phase and frequency flicker noises can be neglected for this purpose. Indeed, the characterization of our Rubidium clock in Figure 6 showed that the frequency flicker noise had a negligible impact on the OASD. Furthermore, the phase white and flicker noises have a similar impact on the standard OASD and cannot be distinguished here. We chose to ignore the phase flicker noise as it is less straightforward to simulate and it should not impact the long term random walk drift that we want to correct. GPS Time can be modeled as pure phase white noise. The corresponding OASD as a function of the averaging time $\tau$ can be modeled [28, 29, 30] by:
$OASD(\tau)\cong A_{WNp}\times\tau^{-1}+A_{WNf}\times\tau^{-1/2}+A_{RWf}\times\tau^{+1/2}.$ (3)
The amplitudes $A$ of these main frequency and phase noises were determined through fitting this model (Eq. 3) to the OASD of the data when characterizing our equipment (see Figure 6) and found to be:
$A_{WNf}$ $=$ $7\times 10^{-12}~{}s^{1/2},$ (4)
@@
for the free-running Rubidium clock and for GPS Time:
$A_{WNf}$ $=$ $0~{}s^{1/2},$ (5)
@@
In Figure 7, the uncorrected simulated signals of the GPS and the clock are reported in dotted symbols for comparison. The increase of the clock’s OASD after $\tau=10^{4}$ s due to the random walk is clearly visible. One can see that with the OASD of the signals (starred symbols) that the correction do eliminate the random walk at longer terms which indicates a success of the method (quadratic). Moreover, one can determine that the ideal length $\Delta t$ of the correction time windows lies around $3\times 10^{4}$ s which corresponds logically to the intersection of the free-running Rubidium clock and GPS Time OASD curves. Indeed, the red curve with a time window of 28800 s shows an ideal combination of the short-term stability of the clock and the absence of random walk at longer scales. On the opposite, the yellow (shorter time window) and light blue (longer time window) curves show respectively a degradation of the short term performance and a remaining random walk component in the region between $\tau=10^{4}$ s and the time window length (here 240000 s).
@@
Finally, it is important to note that although this validates the methods for application on data, those are simplified simulations, in particular because only the two noise types are taken into account. As a result, we do expect differences of performance of the correction on real data. It is also possible that the optimal time window for the correction is slightly different for real data because the simulations are not exact representation of data. Two main differences can be noted: the absence of frequency drift and flicker noises in the simulated Rubidium signal and the fact that we assume a perfect signal to compare the Rubidium signal to when evaluating the OASD. Note that the frequency drift induces a quadratic drift of time signals and should therefore be automatically corrected by our correction method.
@@ 2.2.3 Implementation on data
To check the impact of the correction we compare the Rubidium clock signal before and after correction to UTC(OP) that we receive at the laboratory via the T-REFIMEVE network. The UTC(OP) time signal plays the role of the perfect signal used for the simulations. In the following, we will quantify the stability of the Rubidium signal using the OASD of a time series (according to equation (10) of [31]) consisting of time differences between this signal and UTC(OP). Measuring this time difference frequently, once per second for instance, will allow to also evaluate the very short term stability of the corrected signal which is not possible with the Septentrio measurements that are integrated over $16$ minutes. We use the counter to provide such a measurement every $1$ second approximately. We then perform a simultaneous correction of the Rubidium - GPS Time, as measured by the Septentrio receiver, and of this measured time series. Comparing the OASD of the corrected time series to the uncorrected one, one can quantify the short term stability (below $16$ minutes) after correction while making sure that the random walk was corrected. We can also use this comparison to optimize the value of $\Delta t$ in order to achieve the lowest Allan Standard Deviation possible at all averaging time.
@@ 3 Results
In this Section, we present the results of the correction of the Rubidium clock time stamps obtained for simultaneous measurements of around $35$ days with the Septentrio receiver and the counter. The data used for this analysis are available at this public page. The OASD of the time series measured by the counter is shown in Figure 6. Note that the statistical uncertainty on the estimated OASD, due to the limited number of samples per averaging time, are included as error bars for both curves (Rb and GPS) but they are too small to be visible. Indeed for the Rb vs UTC(OP) OASD, the statistical uncertainty is at the permil level. Up to an averaging time of around $4\times 10^{3}$ s, the stability is limited only by the phase white noise and then by the frequency white noise. After that, the OASD first increases as $\tau^{1/2}$ which is characteristic of the frequency random walk. From $\tau\approx 7\times 10^{4}$ s, the OASD increases proportionally to $\tau$. This is characteristic of a deterministic frequency drift which can be easily characterized and corrected for contrary to the frequency random walk. In comparison, the OASD of the difference between GPS Time and the UTC(OP) reference PPS signal that we receive from LNE-SYRTE, is only limited by a phase white noise at least up to an averaging time of $5\times 10^{5}$ s: the OASD keeps decreasing with the averaging time. At low averaging times, the GPS stability is worse than that of the Rb because of this phase white noise: the GPS OASD is of around $3\times 10^{-12}$ at $960$ s compared to around $2\times 10^{-13}$ OASD for the Rubidium clock. However, at around $10^{4}$ s, the stability of the Rb signal becomes worse compared to GPS Time because of the frequency random walk and drift of the Rubidium clock.
@@
In this paper, we used only the GPS satellites with an elevation angle (angle between line of sight and horizontal direction) larger than $15^{\circ}$ to extract the Rubidium time residuals distribution. During the whole data-taking period, for each data point, the Septentrio receiver was able to track an average of $6.5$ GPS satellites and at least $4$ GPS satellites for each data point. To obtain the Rubidium vs GPS Time difference, we take the mean value of the differences between the Rubidium clock and each GPS satellite tracked in the same integration time window of the Septentrio receiver. The obtained time difference is shown in Figure 10.
@@
Figure 10: Time difference between the Rubidium clock and GPS Time as measured by the Septentrio receiver. The long term quadratic drift is due to the linear frequency drift of the clock. The zoom on the first five days of data also shows shorter term fluctuations caused by the frequency random walk of the Rubidium clock.
It shows that the Rubidium clock time signal drifts away from GPS Time in a quadratic function of time because of the frequency linear drift. After around $35$ days, the difference surpasses $25$ $\mu$s. A zoom on the first five days of data also shows some shorter term fluctuations characteristic of the frequency random walk. Because of those two sources of frequency variations, we see that after a few days of data-taking, the Rubidium clock time signal can drift away from GPS Time by more than a hundred nanoseconds.
@@ 3.1 Offline correction
Figure 11 shows the Allan Standard Deviation of the Rubidium-UTC(OP) data. Note that the measurement rate of the counter was of around $0.995$ measurement per second. The blue curve shows the result for the raw series, before any correction. The other colored curves show the results for the series corrected offline, with different width of the correction time window. Here, we use quadratic fits of the Septentrio data (so $a_{k}\neq 0$ a priori). The shortest time window ($2880$ s) corresponds to approximately $3$ Septentrio $16$ minutes epochs. The medium ($10560$ s) and largest ($240,000$ s) correspond respectively to $11$ and $250$ Septentrio data points.
@@
Figure 11: Offline Correction: Overlapping Allan Standard Deviation of the Rb - UTC(OP) time series before correction (in blue) and after the offline correction with a correction time window of $2880$ s (orange), $10560$ s (green) and $240,000$ s (red). The best stability at both short and long averaging times is obtained for the medium time window ($10560$ s3absent3$\approx$ 3≈ 3 hours).
@@
One sees that with the medium time window compared to the two others, we obtain the best stability at all averaging times. At lower averaging times, the performance is very similar to the uncorrected time series. At higher averaging times, the Allan Standard Deviation is much better than the uncorrected series as it keeps decreasing with increasing $\tau$. This is also the case for correction with the shortest time window. This illustrates the fact that both the $2880$ s and $10560$ s windows are able to correct the frequency random walk and linear drift of the uncorrected time series. However, with the shortest correction time window, the short term stability of the time series is degraded compared to the uncorrected series: the value of the ASD at $100$ s increases by a factor $3$. In this scenario, the corrected Rubidium time signal gets very close to GPS Time which is known to have a higher phase White Noise. Finally, the longest correction time window leads to a similar stability as the shortest one for a small $\tau$, and poorer stability at large $\tau$ (above $5\times 10^{3}$ s).
Figure 12 shows the Rubidium vs GPS Time difference after the offline correction.
@@
Figure 12: Time difference between the Rubidium clock and GPS Time after the offline correction. Three different correction time windows have been tested: $2800$ s (orange), $10560$ s (green) and $240,000$ s (red). These residuals can be compared to the time difference before correction that were shown in Figure 10.
In offline mode, the shorter the correction time window, the lower the residual differences. However, with the medium length time window, we still get time residuals lower than $3$ ns over the whole data-taking period, which is well below the requirements of HK. With the longest correction time window, jumps of a few tens of nanoseconds are introduced in the time residuals. This explains the overall higher ASD: the stability of the signal is limited by those jumps. The time scale of the variations in the data to fit is too small compared to the $240,000$ s time window. In consequence, the fitted tendency from one piece to another is very different, and the fitted piece-wise polynomial is not continuous. It is also interesting, as a cross-check, to have a look at the fluctuations in the time difference between the Rubidium clock and UTC(OP) after correction. This is summarized in the first line of Table 2 that gives the standard deviation of the time series after correction. The deviations with the two shorter correction time windows are indeed very small (below $2$ ns) confirming that this method can be used for synchronization to UTC.
@@
With the offline version of the corrections, we thus obtain a very good synchronization to GPS Time at the level of a few nanoseconds with the $10560$ s time window. However, this version of the correction cannot be applied in real time. In the following, we show the results for the online version of the correction that can be applied in real time to correct the time stamps of events in physics experiments.
@@ 3.2 Online correction
Figure 13 shows the Allan Standard Deviation of the uncorrected (blue) and online corrected (other colors) Rubidium - UTC(OP) times series. The same three correction time windows intervals as before are considered. The top panel shows the results using quadratic fits of the Septentrio data and the bottom panel shows the results with linear fits. For the shortest and medium correction time windows, the linear fits lead to better performance with a lower OASD at low averaging times. At $1000$ s, the OASD with the shortest (medium) correction time window is reduced by a factor $2$ to $3$ (resp. $1.5$).
@@
Figure 13: Online Correction: Overlapping Allan Standard Deviation of the Rb - UTC(OP) time series before correction (in blue) and after the online correction with a correction time window of $2880$ s (orange), $10560$ s (green) and $240,000$ s (red). The data were fitted with quadratic (top) or linear (bottom) functions of time. A better stability, similar to the offline correction, can be obtained using linear fits.
@@
This behavior can be understood by looking at the number of degrees of freedom (number of data points - number of free parameters) in our fits. For the shortest time windows, the number of degrees of freedom is relatively low (00 and $8$) in case of quadratic fits so we risk over-fitting to the past data in order to correct the present data. This number of degrees of freedom is less relevant in the offline correction as the fit is performed on the same data as the correction (the over-fitting is not a problem here). Lowering the number of free parameters is one way of increasing the degrees of freedom hence allowing the fit to better generalize to the present data. Another way to increase the number of degrees of freedom is to increase the number of data points in the fit. For the longest time window, there are $247$ degrees of freedom in the quadratic fit so we lower the risk of over-fitting. On the contrary, in that case, quadratic fits lead to a slightly better correction of the random walk that limits the stability only up to $\tau\sim 3\times 10^{4}$ s whereas with linear fits, it limits the stability up to $\sim 10^{5}$ s. Note that, especially for the shortest correction time window we see a clear degradation of the stability for averaging times lower than the correction window’s length. This is a known effect from linear servo loop theories and periodic perturbations of oscillators [32] and it could be attenuated by scaling down the correction: instead of subtracting the result of the fit, we could subtract only a fraction of it.
@@
Regarding the stability of the corrected Rubidium clock, using linear fits, the conclusions are the same as for the offline correction. The lowest Allan Standard Deviation, for all averaging times, is achieved with the medium width correction time window. With the shortest time window, the short term stability is degraded, and with the longest correction time window, we find poorer long term stability compared to the other corrected scenarios.
@@
If the correction time window is too wide, we cannot correct as well the frequency random walk of the free-running Rubidium: the risk is that the Rubidium time signal locally drifts too far away from GPS Time. This can be observed in the corrected Rubidium against GPS Time in Figure 14 where the maximum difference reaches around $80$ ns (or $25$ ns with quadratic fits) with the $240,000$ s correction time window. With the $10560$ s correction time window, the differences stay in the ±$\pm 5$ ns range. The standard deviation of the time difference with UTC(OP) is also shown in Table 2 for both online corrections. Once again, one can see the reduction of the white noise when using linear instead of quadratic fits. Before correction, as the reader saw in Figure 10, the free-running Rubidium clock can drift away from GPS Time by around $100$ ns in less than $3$ days which means that HK’s requirement for the synchronization with UTC is not met. After online correction with the longest time window tested, the corrected Rubidium time stamps drift by around $60$ ns in a few days because of remaining random walk noise. Even though during the $35$ days data-taking period the time residuals with respect to GPS Time does not exceed $100$ ns, it is not possible to safely claim that the Rubidium clock drift will not exceed HK’s requirement of $100$ ns if we use the $240,000$ s correction time window, because of the random nature of this drift. With shorter time windows, no residual drift is observed, and the residuals are thus contained in a range of a few nanoseconds.
correction time window $2880$ s $10560$ s $240000$ s
offline correction $1.87$ ns $1.79$ ns $5.13$ ns
online correction (quadratic fits) $2.01$ ns $1.83$ ns $9.35$ ns
online correction (linear fits) $1.84$ ns $1.81$ ns $22.66$ ns
Table 2: Standard deviation of the time difference between the Rubidium clock PPS signal and UTC(OP) after correction.